# Flavor diagonal nucleon charges using clover fermions on MILC HISQ ensembles

### v2

@@ III Strategies to Remove Excited-State Contributions
Figure 2: Data for the bare charges $g_{A,T,S}^{u,d}$ (sum of the connected and disconnected contributions) obtained at various separations $\{\tau,t\}$ are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. We show both the “standard” and “$N\pi$” fits, defined in the text, when ESC are manifest. In each case, the result of the fit is shown by lines of the same color as the data for the various $\tau/a$ listed in the label, and the $\tau\to\infty$ value is given by the gray band. We also show, for comparison, the light quark disconnected contributions, $g_{A,S,T}^{l}$, in the panels on the right and the ES fits to them.
Figure 3: Data for the bare strange charges $g_{A,T,S}^{s}$, obtained at various separations $\{\tau,t\}$, are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. These have only disconnected contributions. Only the “standard” fits to remove ESC are shown and considered since the expected lowest multihadron state, $\Sigma K$, is heavier than the nucleon’s radial excitation. Rest is the same as in Fig. 2.
Furthermore, previous studies indicate that the difference in nucleon isovector axial and tensor charges, $g_{A,T}^{u-d}$, from the two strategies is small [11]. This is also observed for $g_{A,T}^{u,d,s}$ in this study, and the errors in results with the “$N\pi$” strategy are comparable to the differences. Only for $g_{S}^{u,d}$ do the differences stand out. Thus, to nail down the best analysis strategy with up to 3-state fits requires future higher statistics studies on more ensembles with $M_{\pi}\lesssim 200$ MeV.
To illustrate the difference in the values of $M_{1}$ coming out of the “standard” and “$N\pi$” analyses, we quote values for the $a09m130$ ensemble on which the difference is the maximum. The Bayesian prior and width used for the mass gap $a\Delta M_{1}=a(M_{1}-M_{0})$ is $0.35(20)$ for the “standard” and $0.12(2)$ for the “$N\pi$” analysis with the central values corresponding to 0.79 GeV and 0.27 GeV in physical units, respectively. While the output $a\Delta M_{1}$ from the simultaneous fits to $C^{\text{2pt}}$ and $C^{\text{3pt}}_{\Gamma}$ could depend on the quark flavor and/or the operator, in practice we get results consistent with the input priors, i.e., in the range $0.24\sim 0.30$ for the “standard” and $0.11\sim 0.14$ for the “$N\pi$” analysis.
@@ IV Model average of various fits made to remove ESC
We have carried out the following twelve ESC fits for both the “standard” and the “$N\pi$” strategies to quantify change under variations of the fit parameters, i.e., the “models”.
  • The full list of $\tau$ values analyzed (covering 0.9–1.5fm) is given in Table 2. The ESC analysis is done (i) with and (ii) without using the data with the largest $\tau$, which have the largest statistical but smallest systematic errors.
  • For each $\tau$, we neglect data on $t_{\text{skip}}$ time slices next to the source and the sink that have the largest ESC. Fits with three different $t_{\text{skip}}$ values in the range of 0.2–0.4 fm were made.
@@
For the final results, we take the average weighted by the AIC score times the inverse of the variance, $\propto\text{exp}[-(\chi^{2}-2N_{\text{dof}})/2]/\sigma^{2}$ [13, 37]. In most cases, one set of $\tau$ and $t_{\text{skip}}$ values gives scores that are much higher than all others for both the 2- and 3-state fit. Thus, in practice, the model average is mainly over the variation of the result with the order of the truncation of Eq. (5). While most 2- and 3-state fits are similar, an example in the “$N\pi$” analysis where we note a qualitative difference, even though the AIC scores were similar, is shown in Fig. 4. Since we do not expect an enhanced $N\pi$ state contribution to $g_{A,T}^{q}$, the large extrapolation in the 2-state fit indicates that this is an artifact due to inadequate statistics. The large error and smaller extrapolation in the 3-state fit is further indication that the data, in this and such cases, are statistics limited. More on this point in Sec. V.
Figure 4: The two panels shows the 2-state (top) and 3-state (bottom) “$N\pi$” fit to the same data for $g_{A}^{d}$ from the $a09m130$ ensemble. Points with $\tau=\{10,12,14,16\}$ and $t_{\text{skip}}=2$ are included in the fits and are shown using filled circles. These two fits have the largest and similar AIC scores and dominate the model average, but differ qualitatively—the 2-state fit shows a large extrapolation, while the 3-state fit has a large error—and give different central values.
Results from a 2- versus 3-state truncation of Eq. (5) highlight the tension between balancing systematic and statistical errors, which is accentuated in the “$N\pi$” analysis. The 2-state fit results may have significant systematic error from residual ESC whereas the statistical precision of the data limits the robustness of the 3-state fits. The same tension is also present between the fits with the two sets of $\tau$ values. The $\chi^{2}/dof$ or the AIC score or the $p$ value of the fit, in most cases, does not distinguish between them. We regard cases with large extrapolation in the ESC fits or a large growth in errors or a large difference between 2- and 3-state fits as outliers due to inadequate statistics rather than a physical effect. The differences are, nevertheless, used to assign an appropriate systematic uncertainty in these cases.
@@ V Pattern of Excited-State Contributions to $g_{A,S,T}^{u,d,s}$
The $a09m130$ ensemble data for the ratio $C^{\text{3pt}}(t;\tau)/C^{\text{2pt}}(\tau)$ are shown in Figs. 2 and  3 as a function of the values of separations $\tau$ and current insertion time $t$ simulated. As already stated above, we choose this physical pion mass ensemble to point out the notable features since the differences in $\Delta M_{1}$ between the “standard” and the “$N\pi$” analyses on it are the largest. In each panel, the final result, the AIC weighted average of the twelve model ESC fits, is shown by the horizontal grey band. The uncertainty bands on data points with a given color are obtained as follows: The value at each $\{\tau,t\}$ point is the weighted average of the twelve model values using the individual AIC weights, and their spread gives the errors. These points are then simply joined to form the bands. Note that these bands are not related in any simple way to the final grey band, but are presented to show the variation in the twelve values, which is given by the thickness of the bands.
@@
  • $g_{A}^{u}$: The convergence of the connected contribution is from below. The disconnected contribution is negative ($\sim-0.06$) and converges to a more negative value as illustrated by $g_{A}^{l,s}$ in Figs. 2 and 3. A notable consequence is that the ESC mostly cancel in the sum and the data at different $\tau$ overlap within the current statistical precision as illustrated for $g_{A}^{u}$ in the top left panel in Fig. 2. We, therefore, estimate the $\tau\to\infty$ value by the weighted average of the data for a range of $\tau$ and $t$ values, i.e., the “plateau method”. For this choice, the “standard” and the “$N\pi$” results are the same.
  • $g_{A}^{d}$: The connected and disconnected contributions are both negative and their ESC add in contrast to what is observed for $g_{A}^{u}$. We, therefore, perform 2- and 3-state fits to obtain the $\tau\to\infty$ value and the extrapolation is significantly larger in the “$N\pi$” analysis.
@@
  • $g_{T}^{u}$ and $g_{T}^{d}$: The ESC contribution to $g_{T}^{u}$ ($g_{T}^{d}$) is significant (small). The disconnected contribution is small ($<0.01$) and negative. Thus, it only marginally reduces $g_{T}^{u}$ and increases $|g_{T}^{d}|$. More important, the fits to $g_{T}^{u}$ and $g_{T}^{d}$ are good and the the size of the extrapolation to the $\tau\to\infty$ value is small.
@@
  • $g_{T}^{l}$ and $g_{T}^{s}$: The values are much smaller than the errors and there is no clear pattern of convergence. We, therefore, assume that there are no significant ESC and make two fits to a constant ansatz—with and without using the full covariance matrix. The results are consistent, however, the fit with the covariance matrix gives a larger error, which we choose to use in the final result.
@@
  • $g_{S}^{u}$ and $g_{S}^{d}$: The connected and disconnected contributions to both are positive, the ESC are large, and the convergence is from below. The size of the extrapolation in the “$N\pi$” fits is large, giving larger values for $g_{S}^{u,d}$, which implies a larger pion-nucleon sigma term [4].
  • $g_{S}^{l}$ and $g_{S}^{s}$: The value is positive and significant, however, the errors in the current data are large. In all but one case, the fits are done with the full covariance matrix for $C^{\text{2pt}}(\tau)$ and $C^{\text{3pt}}_{\Gamma}(t,\tau)$. For $g_{S}^{s}$ from $a09m310$ only, fits with the diagonal approximation to the covariance matrix work better compared to those with the full covariance matrix.
@@
Results for the bare charges extracted using the two strategies to remove ESC, and after model averaging, are summarized in Tables 35 and 4. Their renormalization values in the ${{\overline{\text{MS}}}}$ scheme at the scale $2$ GeV are obtained using the analysis for the 2+1-flavor theory described in Appendix B.
The main question, in addition to the truncation of Eq. (5) to 2- and 3-state fits, is which analysis should be selected for the final values. In the “$N\pi$” analysis, $M_{1}^{N\pi}$ is set to the non-interacting energy of the $N\pi$ state through a narrow prior and is not the output of a data driven fit. While the $N\pi$ state is expected to contribute in principle to all the charges (there is a contribution at one-loop in $\chi$PT  [38], no enhanced contribution is expected except to the scalar charges [4] as evident from Fig. 2. Furthermore, ESC from multihadron states, relative to those from the radial excitations, are expected to become significant only for $M_{\pi}<200$ MeV, i.e., below which the $M_{1}^{N\pi}$ starts to become smaller than $M_{1}^{\rm radial}$. Thus, to distinguish between the “standard” and “$N\pi$” fits, we must rely on data close to the physical pion mass, i.e., in this work on only the $a09m130$ data shown in Figs. 2 and  3. With these caveats, our reasons for choosing the analysis strategy for presenting the final results for each of the charges are given below.
@@
The error in the $\tau\to\infty$ extrapolated value for $g_{A}^{u,d}$ in the “$N\pi$” analysis is large. For $g_{A}^{u}$, we see no significant ESC, so the “standard” fit value is taken. In $g_{A}^{d}$, the ESC to the connected and disconnected terms add and the “$N\pi$” fit value is about $30\%$ larger. We do not have a theoretical reason for such a large effect, nevertheless we take the average of the two estimates. Note, results for the isovector axial charge $g_{A}^{u-d}$, which has no disconnected contributions, are consistent with our final values already given in Ref. [12] as shown in Table 9.
In the case of $g_{T}^{u,d}$, the “standard” and “$N\pi$” analyses give essentially the same central value and the extrapolation is small. Also, the ESC to the connected and disconnected parts mostly cancel in $g_{T}^{d}$. The errors, however, are much larger, beyond what data would indicate, with the “$N\pi$” fit. Since there is no phenomenological or $\chi$PT motivation for a large contribution from multihadron states in the tensor channel, we choose the “standard” analysis.
@@
The data for $g_{S}^{u,d}$ are the cleanest and the extrapolated values have reasonable errors. The two strategies, however, give different values but are not distinguished by the $\chi^{2}/dof$. The size of the extrapolation is large in the “$N\pi$” analysis and, on its own, calls into question even a 3-state fit. On the other hand, $\chi$PT suggests such a enhanced contribution from the $N\pi$ and $N\pi\pi$ states as discussed in Ref. [4]. There we presented results for the pion-nucleon sigma term, $\sigma_{\pi N}$, and highlighted the $\approx 50\%$ difference in $g_{S}^{u,d}$ from the two analysis strategies. The same reasoning leads us to choose the “$N\pi$” analysis in this work. Overall, our conclusion is that $g_{S}^{u,d}$ show large enhanced contributions from the $N\pi$ excited states unlike $g_{A,T}^{u,d}$. For $g_{S}^{s}$, the lowest multihadron state would again be the heavier $\Sigma K$, so no enhancement is expected with respect to the “standard” analysis.
Lastly, even though we have chosen the “standard” analysis values for $g_{A,S}^{s}$, we do show the CCFV extrapolation for the “$N\pi$” analysis data in Figs. 10 and 16. The procedure for the renormalization of these charges is discussed in Appendix B.
@@
Table 3: Final bare axial charges $g_{A}^{u,d,s}$. These $g_{A}^{u-d}$ results are obtained by doing the ES fit to the difference of the connected data with the insertion of the $u$ and $d$ quarks to take into account the correlations between the two. The $g_{A}^{u}$ and $g_{A}^{d}$ are extracted from a fit to the sum of the connected plus disconnected correlation functions. Their sum, $g_{A}^{u}+g_{A}^{d}$, can be compared to the result of a fit to the sum of the correlation functions, $g_{A}^{u+d}$. The errors are calculated using a single elimination jackknife method, and each number is the AIC weighted average over the various model values. The results from the “standard” and “$N\pi$” ESC fits are given separately.
@@ VI Chiral-Continuum-Finite-volume (CCFV) extrapolation and error budget
For the scalar charges $g_{S}^{u}$ and $g_{S}^{d}$, chiral perturbation theory provides two differences. First, the chiral behavior, $g_{S}^{u,d}=d_{0}+d_{a}a+d_{1}M_{\pi}+d_{2}M_{\pi}^{2}+M_{\pi}^{2}\log M_{\pi}^{2}+\ldots$, starts with a term proportional to $M_{\pi}$ [39]. Second, the contribution of $N\pi$ and $N\pi\pi$ excited states from higher order terms (chiral logs, etc.) is large as discussed in [4]. The difference on including these in ESC fits is shown in Fig. 2. Thus, in addition to resolving possibly large ESC, one needs to include the higher order terms (chiral logs, etc.) in the chiral part of the CCFV ansatz. Unfortunately, with data at only three values of $M_{\pi}$, only the first two terms, $\propto M_{\pi}$ and $M_{\pi}^{2}$ in addition to lattice spacing dependence, can be kept. Noting these points, the CCFV ansatz used for the scalar charges $g_{S}^{u,d}$ is
@@
Some specific observations on these CCFV fits to obtain the charges are as follows.

### v3

@@ III Strategies to Remove Excited-State Contributions
Figure 2: Data for the bare charges $g_{A,T,S}^{u,d}$ (sum of the connected and disconnected contributions) obtained at various separations $\{\tau,t\}$ are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. We show both the “standard” and $N\pi$” fits, defined in the text, when ESC are manifest. In each case, the result of the fit is shown by lines of the same color as the data for the various $\tau/a$ listed in the label, and the $\tau\to\infty$ value is given by the gray band. We also show, for comparison, the light quark disconnected contributions, $g_{A,S,T}^{l}$, in the panels on the right and the ES fits to them. The $\chi^{2}/dof$ of the fit with the largest AIC score is given within square parenthesis in the top right corner.
Figure 3: Data for the bare strange charges, $g_{A,T,S}^{s}$, obtained at various separations $\{\tau,t\}$, are plotted versus $(t-\tau/2)/a$ for the physical $M_{\pi}$ ensemble $a09m130$. These have only disconnected contributions. Only the “standard” fits to remove ESC are shown and considered since the expected lowest multihadron state, $\Sigma K$, is heavier than the nucleon’s radial excitation. Rest is the same as in Fig. 2.
Furthermore, previous studies indicate that the difference in nucleon isovector axial and tensor charges, $g_{A,T}^{u-d}$, from the two strategies is small [11]. This is also observed for $g_{A,T}^{u,d,s}$ in this study, and the errors in results with the $N\pi$” strategy are comparable to the differences. Only for $g_{S}^{u,d}$ do the differences stand out. Thus, to nail down the best analysis strategy with up to 3-state fits requires future higher statistics studies on more ensembles with $M_{\pi}\lesssim 200$ MeV.
To illustrate the difference in the values of $M_{1}$ coming out of the “standard” and $N\pi$” analyses, we quote the $a09m130$ values, which have the maximum difference. The Bayesian prior and width used for the mass gap $a\Delta M_{1}=a(M_{1}-M_{0})$ is $0.35(20)$ for the “standard” and $0.12(2)$ for the $N\pi$” analysis with the central values corresponding to 0.79 GeV and 0.27 GeV in physical units, respectively. The output $a\Delta M_{1}$ from the simultaneous fits to $C^{\text{2pt}}$ and $C^{\text{3pt}}_{\Gamma}$ are consistent with the input priors, i.e., they lie in the range $0.24\sim 0.30$ for the “standard” and $0.11\sim 0.14$ for the $N\pi$” analysis. We note little dependence of output values on the quark flavor and/or the operator.
@@ IV Model average of various fits made to remove ESC
We have carried out the following twelve ESC fits, i.e., the 12 “models”, for both the “standard” and the $N\pi$” strategies to quantify change under variations of the fit parameters.
  • The full list of $\tau$ values analyzed (covering 0.9–1.5fm) is given in Table 2. The ESC analysis is done (i) with and (ii) without using the data with the largest $\tau$. These largest $\tau$ data have the largest statistical but smallest systematic errors.
  • For each $\tau$, we neglect data on $\mathop{\tau_{\rm skip}}\nolimits$ time slices next to the source and the sink that have the largest ESC. Fits with three different $t_{\text{skip}}$ values in the range of 0.2–0.4 fm were made.
@@
For the final results, we take the average weighted by the AIC score times the inverse of the variance, $\propto\text{exp}[-(\chi^{2}-2N_{\text{dof}})/2]/\sigma^{2}$ [13, 37]. In most cases, one set of values of $\tau$ and $t_{\text{skip}}$ gives scores that are much higher than all others for both the 2- and 3-state fit. Thus, in practice, the model average is mainly over the variation of the result with the order of the truncation of Eq. (5). While most 2- and 3-state fits are similar, an example in the $N\pi$” analysis where we note a qualitative difference, even though the AIC scores were similar, is shown in Fig. 4. Since we do not expect an enhanced $N\pi$ state contribution to $g_{A,T}^{q}$, the large extrapolation in the 2-state fit indicates that this is an artifact due to inadequate statistics. The large error and smaller extrapolation in the 3-state fit is further indication that the data, in this and such cases, are statistics limited. More on this point in Sec. V.
Figure 4: The two panels shows the 2-state (top) and 3-state (bottom) $N\pi$” fit to the same data for $g_{A}^{d}$ from the $a09m130$ ensemble. Points with $\tau=\{10,12,14,16\}$ and $t_{\text{skip}}=2$ are included in the fits and are shown using filled circles. These two fits have the largest and similar AIC scores and dominate the model average, but differ qualitatively—the 2-state fit shows a large extrapolation, while the 3-state fit has a large error—and give different central values.
Results from a 2- versus 3-state truncation of Eq. (5) highlight the tension between balancing systematic and statistical errors, which is accentuated in the $N\pi$” analysis. The 2-state fit results may have significant systematic error from residual ESC whereas the statistical precision of the data limits the robustness of the 3-state fits. The same tension is also present between the fits with the two sets of $\tau$ values. The $\chi^{2}/dof$ or the AIC score or the $p$ value of the fit, in most cases, does not distinguish between them. We regard cases with large extrapolation in the ESC fits or a large growth in errors or a large difference between 2- and 3-state fits as outliers due to inadequate statistics rather than a physical effect. The differences are, nevertheless, used to assign an appropriate systematic uncertainty in these cases.
@@ V Pattern of Excited-State Contributions to $g_{A,S,T}^{u,d,s}$
The $a09m130$ ensemble data for the ratio $C^{\text{3pt}}(t;\tau)/C^{\text{2pt}}(\tau)$ are shown in Figs. 2 and  3 as a function of the values of separations $\tau$ and current insertion time $t$ simulated. As already stated above, we choose this physical pion mass ensemble to point out the notable features since the differences in $\Delta M_{1}$ between the “standard” and the $N\pi$” analyses on it are the largest. In each panel, the final result, the AIC weighted average of the twelve ESC fits, the models, is shown by the horizontal grey band. The uncertainty bands on data points with a given color are obtained as follows: The value at each $\{\tau,t\}$ point is the weighted average of the twelve model values using the individual AIC weights, and the ”error” is the spread. These points are then simply joined to form the bands. Note that these bands are not related in any simple way to the final grey band, but are presented to show the variation in the twelve values, which is given by the thickness of the bands.
@@
  • $g_{A}^{u}$: The convergence of the connected contribution is from below. The disconnected contribution is negative ($\sim-0.06$) and converges to a more negative value as illustrated by $g_{A}^{l,s}$ in Figs. 2 and 3. A notable consequence is that the ESC mostly cancel in the sum and the data at different $\tau$ overlap within the current statistical precision as illustrated for $g_{A}^{u}$ in the top left panel in Fig. 2. We, therefore, estimate the $\tau\to\infty$ value by the weighted average of the data for a range of $\tau$ and $t$ values, i.e., the “plateau method”. Thus, the “standard” and the $N\pi$” results are the same.
  • $g_{A}^{d}$: The connected and disconnected contributions are both negative and their ESC add in contrast to what is observed for $g_{A}^{u}$. We, therefore, perform 2- and 3-state fits to obtain the $\tau\to\infty$ value and the extrapolation is significantly larger in the $N\pi$” analysis.
@@
  • $g_{T}^{u}$ and $g_{T}^{d}$: The ESC contribution to $g_{T}^{u}$ ($g_{T}^{d}$) is significant (small). The disconnected contribution is small (<0.01<0.01) and negative. Thus, it only marginally reduces $g_{T}^{u}$ and increases $|g_{T}^{d}|$. More important, the fits to $g_{T}^{u}$ and $g_{T}^{d}$ are good and the size of the extrapolation to the $\tau\to\infty$ value is small.
@@
  • $g_{T}^{l}$ and $g_{T}^{s}$: The values are much smaller than the errors and there is no clear pattern of convergence. We, therefore, assume that there are no significant ESC and make two fits to a constant ansatz—with and without using the full covariance matrix. The results are consistent, however, the fit with the covariance matrix gives a larger error, and is chosen for the final result.
@@
  • $g_{S}^{u}$ and $g_{S}^{d}$: The connected and disconnected contributions to both are positive, the ESC are large, and the convergence is from below. The size of the extrapolation in the $N\pi$” fits is large, giving larger values for $g_{S}^{u,d}$, which implies a larger pion-nucleon sigma term [4].
  • $g_{S}^{l}$ and $g_{S}^{s}$: Their values are also positive and significant, however, the errors in the current data are large. In one case, $g_{S}^{s}$ from the $a09m310$ ensemble, only the diagonal elements of the covariance matrix were used since we did not find reasonable ranges of $\{\tau,t,t_{\rm skip}\}$ for which the off-diagonal elements were stable.
@@
Results for the bare charges extracted using the two strategies to remove ESC, and after model averaging, are summarized in Tables 35 and 4. Their renormalization and conversion to the $\overline{\text{MS}}$}} scheme at the scale $2$ GeV is done using the analysis for the 2+1-flavor theory described in Appendix B.
[…]
@@
The error in the $\tau\to\infty$ extrapolated value for $g_{A}^{u,d}$ in the $N\pi$” analysis is large. For $g_{A}^{u}$, we see no significant ESC, so the “standard” fit value is taken. In $g_{A}^{d}$, the ESC to the connected and disconnected terms add and the $N\pi$” fit value is about $30\%$ larger. We do not have a theoretical reason for such a large effect, nevertheless we take the average of the two estimates. Note, results for the isovector axial charge $g_{A}^{u-d}$, which has no disconnected contributions, are consistent with our final values already given in Ref. [12] as shown in Table 9.
[…]
@@
[…]
Lastly, even though we have chosen the “standard” analysis values for $g_{A,S}^{s}$, we do show the CCFV extrapolation of the data from the $N\pi$” analysis in Figs. 10 and 16. The procedure for the renormalization of these charges is discussed in Appendix B.
@@
Table 3: Final bare axial charges $g_{A}^{u,d,s}$. These $g_{A}^{u-d}$ results are obtained by doing the ES fit to the difference of the connected data with the insertion of the $u$ and $d$ quarks to take into account the correlations between the two. The $g_{A}^{u}$ and $g_{A}^{d}$ are extracted from a fit to the sum of the connected plus disconnected correlation functions. Their sum, $g_{A}^{u}+g_{A}^{d}$, can be compared to the result of a fit to the sum of the correlation functions, $g_{A}^{u+d}$. The errors are calculated using a single elimination jackknife method using data binned as described in the text, and each number is the AIC weighted average over the various model values. The results from the “standard” and $N\pi$” ESC fits are given separately.
@@ VI Chiral-Continuum-Finite-volume (CCFV) extrapolation and error budget
For the scalar charges $g_{S}^{u}$ and $g_{S}^{d}$, chiral perturbation theory gives two differences. First, the chiral behavior, $g_{S}^{u,d}=d_{0}+d_{a}a+d_{1}M_{\pi}+d_{2}M_{\pi}^{2}+M_{\pi}^{2}\log M_{\pi}^{2}+\ldots$, starts with a term proportional to $M_{\pi}$ [39]. Second, the contribution of $N\pi$ and $N\pi\pi$ excited states from higher order terms (chiral logs, etc.) is large as discussed in [4]. The difference on including these in ESC fits is shown in Fig. 2. Thus, one needs to include the higher order terms (chiral logs, etc.) in both, fits to remove possibly large ESC and in the chiral part of the CCFV ansatz. Unfortunately, with data at only three values of $M_{\pi}$, only the first two terms, $\propto M_{\pi}$ and $M_{\pi}^{2}$ in addition to lattice spacing dependence, can be kept without overparameterization. Noting these points, the CCFV ansatz used for the scalar charges $g_{S}^{u,d}$ is
@@
Some specific observations on these CCFV fits to obtain the charges at the physical point are as follows.